# On the Friendly number of 10

### v1

@@ Abstract.
To answer the question ”Is $10$ a solitary number?”, we have established a necessary and sufficient condition on primes $p$ and $q$ so that $p$ divides $\sigma(q^{2a})$($a\in\mathbb{Z^{+}}$). Also We have found the form of $q$ whenever $p$ divides $\sigma(q^{2a})$ for $q>p$ Using these two, we have proved that if $n$ is a friend of $10$ then it is a odd square with at least seven distinct prime factors, $5$ being the least one. Moreover there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$, provided $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Further, there exist prime factors $p$ and $q$ not necessarily distinct of $n$ such that $p\equiv 1\pmod{10}$ and $q\equiv 1\pmod{6}$. Besides, we have proved that if a Fermat prime $F_{k}$ divides $n$ then $n$ must have a prime factor congruent to $1$ modulo $2F_{k}$.
footnotetext: Mathematics Subject Classification (2020):11A25.footnotetext: Keywords: Abundancy Index, Sum of Divisors,Friendly Numbers, Solitary Numbers.
@@ 1. Introduction
For any positive integer $n$, $\sigma(n)$ denotes the sum of the positive divisors of $n$. The abundancy index of $n$ is defined as $\frac{\sigma(n)}{n}$. We use $I(n)=\frac{\sigma(n)}{n}$ to denote the abundancy index throughout in this paper. A positive integer $n$ is said to be friendly number if there exist a positive integer $m$ other than $n$ such that it has same abundancy index as $n$ i.e., $I(n)=I(m)$ and in that case we say $m$ is a friend of $n$ and if $n$ has no friend then $n$ is called a solitary number. In [2] we can find that if $(\sigma(n),n)=1$ then $n$ is a solitary number. For a prime number $p$ we have $\sigma(p)=p+1$ thus $(\sigma(p),p)=1$ which shows that any prime $p$ is a solitary number. Further using this we can show that all powers of primes are in fact solitary. $10$ is the smallest positive integer whose classification in terms of solitary and friendly is unknown as in case of $1,2,3,4,5,7,8,9$ we can use $(\sigma(n),n)=1$ this fact to prove them solitary and for $6$ it is well known that it has many friends $28,496,8128$ etc. (all friends of $6$ are called perfect numbers). Although all friends of $6$ are found even and therefore one question naturally raised that ”are there any odd perfect numbers?” which is still an open problem in mathematics. However in this paper we will discuss on the friend of $10$ mainly and we will prove the following theorems.
@@
Let $n$ be a friend of $10$ with $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Then there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$.
@@
If $p$ and $q$ are two distinct prime numbers with $q>p$, then $p$ divides $\sigma(q^{2a}),\leavevmode\nobreak\ a\in\mathbb{Z^{+}}$ if and only if for $r\neq 1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ such that $r^{f}\equiv 1\pmod{p}$ and for $r=1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{p}$.
@@ 2. Properties of the Abundancy Index
In this section, we will prove some elementary properties of the abundancy index, which will not only help us to understand the behavior of the abundancy index but also help us to prove some properties of the friend of $10$. These properties will play a crucial role in proving our main theorems.
@@
Proof.
Let $a$ and $b$ be two positive integers such that $(a,b)=1$. Then $\sigma(ab)=\sigma(a)\sigma(b)$ as $\sigma(n)$ is a weakly multiplicative function and therefore $I(ab)=\frac{\sigma(ab)}{ab}=\frac{\sigma(a)\sigma(b)}{ab}=\frac{\sigma(a)}{a}\frac{\sigma(b)}{b}=I(a)I(b)$. This proves that $I(n)$ is weakly multiplicative. ∎
@@
Proof.
Let $a,n$ be two positive integers with $a>1$. Then
$\sigma(an)$ $=\sum_{d\mid an}d=\sum_{d^{{}^{\prime}}\mid n}ad^{{}^{\prime}}+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}$
$=a\sum_{d^{{}^{\prime}}\mid n}d^{{}^{\prime}}+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}$
$=a\sigma(n)+\sum_{{\begin{subarray}{c}d^{{}^{\prime\prime}}\mid an\\ d^{{}^{\prime\prime}}\neq ad^{{}^{\prime}}\\ \text{where\leavevmode\nobreak\ }d^{{}^{\prime}}\mid n\end{subarray}}}d^{{}^{\prime\prime}}.$
Therefore $\frac{\sigma(an)}{a\sigma(n)}>1$ and this proves $I(an)>I(n)$. ∎
@@
Proof.
Since
$\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}=\prod_{i=1}^{m}\sigma(p_{i}^{a_{i}})=\prod_{i=1}^{m}\biggl{(}\sum_{j=0}^{a_{i}}p_{i}^{j}\biggl{)}=\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}-1}$
We have
$I\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}=\frac{\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{a_{i}}\biggl{)}}{\prod_{i=1}^{m}p_{i}^{a_{i}}}=\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}=\prod_{i=1}^{m}\biggl{(}\sum_{j=0}^{a_{i}}p_{i}^{-j}\biggl{)}$
This completes the proof. ∎
@@
Proof.
For $1\leq i\leq m$, let us define $f_{i}:(1,\infty)\longrightarrow\mathbb{R}$ by $f_{i}(x)=\frac{x^{t_{i}+1}-1}{x^{t_{i}+1}-x^{t_{i}}}$, for some positive integer $t_{i}$. Then
$f_{i}^{{}^{\prime}}(x)=\frac{t_{i}x^{t_{i}-1}(x-1)-x^{t_{i}}(x^{t_{i}}-1)}{(x^{t_{i}+1}-x^{t_{i}})^{2}}$
$=\frac{x^{t_{i}-1}(x-1)\{t_{i}-x(x^{t_{i}-1}+x^{t_{i}-2}+...+1)\}}{(x^{t_{i}+1}-x^{t_{i}})^{2}}.$
Clearly, as $x>1,t_{i}\geq 1$, $t_{i}x^{t_{i}-1}(x-1)>0$ and $(x^{t_{i}+1}-x^{t_{i}})^{2}>0$ but $t_{i}-x(x^{t_{i}-1}+x^{t_{i}-2}+...+1)\}<0$, therefore we have $f_{i}^{{}^{\prime}}(x)<0$ for all $x\in(1,\infty)$. This shows that $f_{i}$ is a decreasing function for each $x\in(1,\infty)$ and for all $1\leq i\leq m$. Now let $p_{i}$, $q_{i}\in(1,\infty)$ such that $p_{i}\leq q_{i}$, for all $1\leq i\leq m$. Then for all $1\leq i\leq m$, $f_{i}(p_{i})\geq f_{i}(q_{i})$ i.e.,
$\frac{p_{i}^{t_{i}+1}-1}{p_{i}^{t_{i}}(p_{i}-1)}\geq\frac{q_{i}^{t_{i}+1}-1}{q_{i}^{t_{i}}(p_{i}-1)}$
Therefore
$\prod_{i=1}^{m}\frac{p_{i}^{t_{i}+1}-1}{p_{i}^{t_{i}}(p_{i}-1)}\geq\prod_{i=1}^{m}\frac{q_{i}^{t_{i}+1}-1}{q_{i}^{t_{i}}(p_{i}-1)}$
implies
$\frac{\sigma\biggl{(}\prod_{i=1}^{m}p_{i}^{t_{i}}\biggl{)}}{\prod_{i=1}^{m}p_{i}^{t_{i}}}\geq\frac{\sigma\biggl{(}\prod_{i=1}^{m}q_{i}^{t_{i}}\biggl{)}}{\prod_{i=1}^{m}q_{i}^{t_{i}}}.$
This proves $I\biggl{(}\prod_{i=1}^{m}p_{i}^{t_{i}}\biggl{)}\geq I\biggl{(}\prod_{i=1}^{m}q_{i}^{t_{i}}\biggl{)}$.
@@
Proof.
Since, for every $1\leq i\leq m$ we have
$\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}<\frac{p_{i}^{a_{i}+1}}{p_{i}^{a_{i}}(p_{i}-1)}=\frac{p_{i}}{p_{i}-1}$
Therefore
$\prod_{i=1}^{m}\frac{p_{i}^{a_{i}+1}-1}{p_{i}^{a_{i}}(p_{i}-1)}<\prod_{i=1}^{m}\frac{p_{i}}{p_{i}-1}$
i.e., $I(n)<\prod_{i=1}^{m}\frac{p_{i}}{p_{i}-1}$ and this proofs the lemma.
@@ 3. Basic Properties of friend of 10
In this section, we will prove some basic properties of the friend of $10$, which will help us to understand the nature of the friend of $10$. Moreover these results will play a significant role in the proofs of our main theorems.
@@ 4. Proofs of the Main Theorems
Conversely, let $2a+1\equiv 0\pmod{f}$ and $q^{f}\equiv 1\pmod{p^{k}}$ where $f$ is the smallest odd positive integer greater than $1$. Then $2a+1=tf$, for some $t\in\mathbb{Z^{+}}$ and therefore
@@
Let $n=\prod_{i=1}^{m}p_{i}^{2a_{i}}$$(p_{1}=5)$ with $5^{2a}\mid\mid n$, ($a\in\mathbb{Z^{+}}$) be a friend of $10$. In particular let $q=5$ in Theorem 1.1. Then there must exists a prime $p$ of $n$ such that $p\mid\sigma(5^{2a})$. Since $p\geq 5$, $p\nmid 4$, therefore we have $2a+1\equiv 0\pmod{f}$ and $5^{f}\equiv 1\pmod{p}$ where $f$ is the smallest odd positive integer greater than 1. As $f\mid(2a+1)$ and $f\mid\phi(p)$ i.e., $f\mid(p-1)$, we have $f\leq\min\{2a+1,p-1\}.$
@@
Let $n$ be a friend of $10$ and $F_{k}=2^{2^{k}}+1$ ($k\in\mathbb{Z}_{\geq 0}$) be a prime factor of $n$. If possible, let $n$ has no prime factor congruent to $1$ modulo $F_{k}$. Now since $F_{k}\mid n$, by Theorem 1.1 there exists a prime factor $p$ of $n$ with $p^{2a_{p}}\mid\mid n$ ($a_{p}\in\mathbb{Z^{+}}$) and $F_{k}^{k^{*}-1}\mid\mid(p-1)$ such that $2a_{p}+1\equiv 0\pmod{f}$ and $p^{f}\equiv 1\pmod{F_{k}^{k^{*}}}$ where $f$ is smallest positive odd integer greater than $1$. Since no prime factors $p$ of $n$ is congruent to $1$ modulo $F_{k}$, we have $k^{*}=1$. Further as $f\mid\phi(F_{k}^{k^{*}})$ i.e., $f\mid\phi(F_{k})$ and $\phi(F_{k})=2^{2^{k}}$ this shows that there does not exist any odd positive divisor of $\phi(F_{k})$ greater than $1$ as it has only one prime factor $2$. Therefore by Theorem 1.1 we can conclude that $F_{k}$ does not divide any $\sigma(p^{2a_{p}})$. Since $n$ is a friend of 10 using Remark 4.1 we can conclude that $F_{k}$ can not be a divisor of $n$ but this contradicts our assumption that $F_{k}$ is a prime factor of $n$. Hence there must exists a prime factor $p$ of $n$ such that $p\equiv 1\pmod{F_{k}}$. Since $(p-1)$ is even we have $p\equiv 1\pmod{2F_{k}}$.∎
@@
(4.1) $\sum_{j=0}^{2a}q^{j}\equiv 0\pmod{p}$
As $q>p$, we can write $q=pk+r$, for some $k\in\mathbb{Z^{+}}$ and $0\leq r\leq(p-1)$. If $r=0$ then $q=pk$ but $p\nmid\sigma((pk)^{2a})$, therefore $1\leq r\leq(p-1)$. From (4.1) we have that
@@
Suppose that $n$ has exactly two distinct odd prime factors $p_{1}=5$ and $p_{2}$ where $p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}$ for some positive integers $a_{1}$, $a_{2}$. By Lemma 2.4 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$ but Lemma 2.5 shows that $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}=\frac{35}{24}>I(5^{2a_{1}}\cdot 7^{2a_{2}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than two distinct prime factors.
Case-2:
@@
Suppose that $n$ has exactly three distinct odd prime factors $p_{1}=5$, $p_{2}$ and $p_{3}$ with $p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}$ where $a_{1}$, $a_{2}$, $a_{3}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}=\frac{77}{48}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than there distinct prime factors.
Case-3:
@@
Suppose that $n$ has exactly four distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$ and $p_{4}$ with $p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{13}{12}=\frac{1001}{576}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than four distinct prime factors.
Case-4:

### v2

@@ Abstract.
To answer the question ”Is $10$ a solitary number?”, we have established a necessary and sufficient condition on primes $p$ and $q$ so that $p$ divides $\sigma(q^{2a})$($a\in\mathbb{Z^{+}}$). Also We have found the form of $q$ whenever $p$ divides $\sigma(q^{2a})$ for $q>p$ Using these two, we have proved that if $n$ is a friend of $10$ then it is an odd square with at least seven distinct prime factors, $5$ being the least one. Moreover there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$ such that $5^{f}\equiv 1\pmod{p}$, provided $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Further, there exist prime factors $p$ and $q$ not necessarily distinct of $n$ such that $p\equiv 1\pmod{10}$ and $q\equiv 1\pmod{6}$. Besides, we have proved that if a Fermat prime $F_{k}$ divides $n$ then $n$ must have a prime factor congruent to $1$ modulo $2F_{k}$.
footnotetext: Keywords: Abundancy Index, Sum of Divisors, Friendly Numbers, Solitary Numbers.footnotetext: MSC2020: 11A25
@@ 1. Introduction
For any positive integer $n$, $\sigma(n)$ denotes the sum of the positive divisors of $n$. The abundancy index of $n$ is defined as $\frac{\sigma(n)}{n}$. We use $I(n)=\frac{\sigma(n)}{n}$ to denote the abundancy index throughout in this paper. A positive integer $n$ is said to be friendly number if there exist a positive integer $m$ other than $n$ such that it has same abundancy index as $n$ i.e., $I(n)=I(m)$ and in that case we say $m$ is a friend of $n$ and if $n$ has no friend then $n$ is called a solitary number. In [2] we can find that if $(\sigma(n),n)=1$ then $n$ is a solitary number. For a prime number $p$ we have $\sigma(p)=p+1$ thus $(\sigma(p),p)=1$ which shows that any prime $p$ is a solitary number. Further using this we can show that all powers of primes are in fact solitary. $10$ is the smallest positive integer whose classification in terms of solitary and friendly is unknown as in case of $1,2,3,4,5,7,8,9$ we can use $(\sigma(n),n)=1$ this fact to prove them solitary and for $6$ it is well known that it has many friends $28,496,8128$ etc. (all friends of $6$ are called perfect numbers). Although all friends of $6$ are found even and therefore one question naturally raised that ”are there any odd perfect numbers?” which is still an open problem in mathematics. J. Ward [6] proved that if 10 has a friend then it is an odd square positive integer having at least six distinct prime divisors with $5$ being the least one and it has at least one prime divisor congruent to $1$ modulo $3$. In this paper we improve his results and we prove the following,
@@
Let $n$ be a friend of $10$ with $5^{2a}\mid\mid n$, $a\in\mathbb{Z^{+}}$. Then there exists a prime factor $p$ of $n$ such that $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ and less than or equal to $\min\{2a+1,p-1\}$ such that $5^{f}\equiv 1\pmod{p}$.
@@
If $p$ and $q$ are two distinct prime numbers with $q>p$, then $p$ divides $\sigma(q^{2a}),~{}a\in\mathbb{Z^{+}}$ if and only if for $r\neq 1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ such that $r^{f}\equiv 1\pmod{p}$ and for $r=1$, $q\equiv r\pmod{p}$ and $2a+1\equiv 0\pmod{p}$.
@@ 2. Properties of the Abundancy Index
In this section, we will note some useful lemmas which will help us to prove our main results.
@@ 3. Basic Properties of friend of 10
In this section, we will prove some basic properties of the friend of $10$, which will help us to understand the nature of the friend of $10$. Moreover these results will play a significant role in proving our main results.
@@ 4. Proofs of the Main Theorems
Conversely, let $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than $1$ such that $q^{f}\equiv 1\pmod{p^{k}}$. Then $2a+1=tf$, for some $t\in\mathbb{Z^{+}}$ and therefore
@@
Let $n=\prod_{i=1}^{m}p_{i}^{2a_{i}}$$(p_{1}=5)$ with $5^{2a}\mid\mid n$, ($a\in\mathbb{Z^{+}}$) be a friend of $10$. In particular let $q=5$ in Theorem 1.1. Then there must exists a prime $p$ of $n$ such that $p\mid\sigma(5^{2a})$. Since $p\geq 5$, $p\nmid 4$, therefore we have $2a+1\equiv 0\pmod{f}$ where $f$ is the smallest odd positive integer greater than 1 such that $5^{f}\equiv 1\pmod{p}$. As $f\mid(2a+1)$ and $f\mid\phi(p)$ i.e., $f\mid(p-1)$, we have $f\leq\min\{2a+1,p-1\}.$
@@
Let $n$ be a friend of $10$ and $F_{k}=2^{2^{k}}+1$ ($k\in\mathbb{Z}_{\geq 0}$) be a prime factor of $n$. If possible, let $n$ has no prime factor congruent to $1$ modulo $F_{k}$. Now since $F_{k}\mid n$, by Theorem 1.1 there exists a prime factor $p$ of $n$ with $p^{2a_{p}}\mid\mid n$ ($a_{p}\in\mathbb{Z^{+}}$) and $F_{k}^{k^{*}-1}\mid\mid(p-1)$ such that $2a_{p}+1\equiv 0\pmod{f}$ where $f$ is smallest positive odd integer greater than $1$ satisfying $p^{f}\equiv 1\pmod{F_{k}^{k^{*}}}$. Since no prime factors $p$ of $n$ is congruent to $1$ modulo $F_{k}$, we have $k^{*}=1$. Further as $f\mid\phi(F_{k}^{k^{*}})$ i.e., $f\mid\phi(F_{k})$ and $\phi(F_{k})=2^{2^{k}}$ this shows that there does not exist any odd positive divisor of $\phi(F_{k})$ greater than $1$ as it has only one prime factor $2$. Therefore by Theorem 1.1 we can conclude that $F_{k}$ does not divide any $\sigma(p^{2a_{p}})$. Since $n$ is a friend of 10 using Remark 4.1 we can conclude that $F_{k}$ can not be a divisor of $n$ but this contradicts our assumption that $F_{k}$ is a prime factor of $n$. Hence there must exists a prime factor $p$ of $n$ such that $p\equiv 1\pmod{F_{k}}$. Since $(p-1)$ is even we have $p\equiv 1\pmod{2F_{k}}$.∎
@@
(1) $\sum_{j=0}^{2a}q^{j}\equiv 0\pmod{p}$
As $q>p$, we can write $q=pk+r$, for some $k\in\mathbb{Z^{+}}$ and $0\leq r\leq(p-1)$. If $r=0$ then $q=pk$ but $p\nmid\sigma((pk)^{2a})$, therefore $1\leq r\leq(p-1)$. From (1) we have that
@@
Case-1: Suppose that $n$ has exactly two distinct odd prime factors $p_{1}=5$ and $p_{2}$ where $p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}$ for some positive integers $a_{1}$, $a_{2}$. By Lemma 2.4 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$ but Lemma 2.5 shows that $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}=\frac{35}{24}>I(5^{2a_{1}}\cdot 7^{2a_{2}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than two distinct prime factors.
@@
Case-2: Suppose that $n$ has exactly three distinct odd prime factors $p_{1}=5$, $p_{2}$ and $p_{3}$ with $p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}$ where $a_{1}$, $a_{2}$, $a_{3}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}=\frac{77}{48}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than there distinct prime factors.
@@
Case-3: Suppose that $n$ has exactly four distinct odd prime factors $p_{1}=5$, $p_{2}$, $p_{3}$ and $p_{4}$ with $p_{4}>p_{3}>p_{2}>p_{1}=5$. Then we can write $n=5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}}$ where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$ are positive integers. Again by Lemma 2.4 and Lemma 2.5 we have $I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})\geq I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$ and $\frac{9}{5}>\frac{5}{4}\cdot\frac{7}{6}\cdot\frac{11}{10}\cdot\frac{13}{12}=\frac{1001}{576}>I(5^{2a_{1}}\cdot 7^{2a_{2}}\cdot 11^{2a_{3}}\cdot 13^{2a_{4}})$ i.e., $\frac{9}{5}>I(5^{2a_{1}}\cdot p_{2}^{2a_{2}}\cdot p_{3}^{2a_{3}}\cdot p_{4}^{2a_{4}})=I(n)$. Therefore $n$ can’t be a friend of $10$ and $n$ has more than four distinct prime factors.